Invoice Report

December 1993

NMPC Project 07-9199

High Performance Computing and Parallel Processing in the Utility Industry

Geoffrey Fox, Paul Coddington, Alok Choudhary, Salim Hariri, Michael Nilan, Sanjay Ranka

Northeast Parallel Architectures Center, Syracuse University

Work completed on this project for the period December 1 to December 31 1993:

• Improvements to a prototype parallel sparse matrix program for power systems problems.

• Research into the requirents for parallelization of the EPRI ETMSP transient stability analysis program, and work on interfacing the program to a parallel sparse matrix solver.

• Study of the problems in interfacing state-of-the-art Differential-Algebraic Equation Solvers into ETMSP.

• Development of a prelinary version of an expert system to analyze transient stability output.

• Development of a preliminary implementation of a simulated annealing optimization program for unit commitment.

• Studies of the basic computational issues relevant to power systems application programs, such as the main algorithms used in programs, methods for parallelizing these algorithms, and the most appropriate computer languages and architectures.

• Further work on a preliminary survey of the applications programs used at NMPC and the applicability of parallel processing to these programs.

• Attendance of 3 EPRI conferences on programs and expert systems for power systems problems of relevance to our project.

See attached financial summary for the expenditures for this month for the work described in this invoice, and for expenditures to-date for this project.

The work described here involved 100% of one month’s work time for 1 researcher, 20% of 1 month’s work time for 1 faculty member, 5% of 1 month’s work time for 2 additional faculty members, 3% of 1 month’s work time for 1 additional faculty member, and 65% of 1 month’s work time for 5 graduate students and 1 software engineer.

Travel expenses were for NPAC representatives to attend 3 EPRI conferences.

Total including fringe benefits and overhead is $32,680.92 for this invoice, $108,860.51 todate.